A food safety guideline is that the mercury in fish should be below 1 part per million​ (ppm). Listed below are the amounts of mercury​ (ppm) found in tuna sushi sampled at different stores in a major city. Construct a 95​% confidence interval estimate of the mean amount of mercury in the population. Does it appear that there is too much mercury in tuna​ sushi? 0.57  0.81  0.11  0.96  1.24  0.58  0.80 What is the confidence interval estimate of the population mean mu​?
Answer #1

Solution:

Solution:

x x2
0.57 0.3249
0.81 0.6561
0.11 0.0121
0.96 0.9216
1.24 1.5376
0.58 0.3364
0.8 0.64
\sumx = 5.07 \sumx2 = 4.4287

Mean \bar x    = (\sumx / n) )The sample mean is \bar x

= (0.57+0.81+0.11+0.96+1.24+0.58+0.80 / 7 )

= 5.07 / 7

= 0.7243

Mean \bar x    = 0.72

c ) The sample standard is S

  S = \sqrt{} ( \sum x2 ) - (( \sum x)2 / n ) n -1

= \sqrt{} (4.4287 ( (5.07 )2 / 7 ) 6

   = \sqrt{} (4.4287 - 3.6721 / 6)

=\sqrt{} (0.7566 / 6 )

= \sqrt{} 0.1261

= 0.3551

The sample standard is = 0.35

Degrees of freedom = df = n - 1 = 7 - 1 = 6

At 95% confidence level the t is ,

\alpha  = 1 - 95% = 1 - 0.95 = 0.05

\alpha / 2 = 0.05 / 2 = 0.025

t\alpha /2,df = t0.025,6 =2.447

Margin of error = E = t\alpha/2,df * (s /\sqrtn)

= 2.447 * (0.35 / \sqrt 7)

= 0.32

Margin of error = 0.32

The 95% confidence interval estimate of the population mean is,

\bar x - E < \mu < \bar x + E

0.73 - 0.32 < \mu < 0.72 + 0.32

0.41 < \mu < 1.04

( 0.41. 104 )